A book is thrown straight upward from the top of the MASH building with an initial speed of 14m/s. If the building is 92m tall, A).how long is the book in the air?B). What velocity will the book hit the ground with? Ignore air resistance.


1
Expert's answer
2021-02-15T17:38:21-0500

a)


"-V=v-gt\\-44.7=14-9.8t\\t=6.0\ s"

b)


"V^2=v^2+2gh\\V^2=14^2+2(9.8)(92)\\V=44.7\frac{m}{s}"